OpenStudy (anonymous):

Slope=-8;(-2,2) a.8x+y=18 b.8x+y=-18 c.-8x+y=-18 d.8x-y=-18

OpenStudy (anonymous):

dont u mean C.

OpenStudy (anonymous):

oh its a cause u change the sings?

OpenStudy (anonymous):

*signs

OpenStudy (anonymous):

point slope form line equation is y -y1 = m (x - x1) y - 2 = -8 (x - (-2) y - 2 = -8 ( x + 2) y - 2 = -8 x - 16 8x +y = - 16 +2 8x +y = - 14

OpenStudy (mathmate):

Can you check if the question wasn't Slope=-8;(-2,-2)

OpenStudy (anonymous):

point slope form line equation is y -y1 = m (x - x1) y - (-2) = -8 (x - (-2) y + 2 = -8 ( x + 2) y + 2 = -8 x - 16 8x +y = - 16 -2 8x +y = - 18

OpenStudy (anonymous):

I calculator wrong y-2=-8(x-(-2)) y= -8x-16+2 y = -8x -14 the answer is none of these